# Juding the Judges: A Systematic Investigation of Position Bias in Pairwise Comparative Assessments by LLMs

### v6

@@ 4 Main Findings
Position Bias is weakly Length-dependent We investigate the impact of three different lengths on the position bias of LLM judges: the length of the question (task input length), the solution length of candidate models (task output length), and the length of the entire prompt (prompt length). Since the tokenization methods vary by models, we apply the length of the prompt component strings for uniformity and scalability. The analysis from Fig. 4(d)(e)(f) shows that there is very weak relationship between the lengths of prompt components and position bias. By stepwise regression, the average task output length is only significant in predicting $PF$, adding a minimal change in AIC as shown in Appendix Table. 4. Therefore, the impact of lengths on position bias is minimal compared to answer quality gap and judge or task level variances.
Practical Implications for Benchmark Design We complement our investigations of position bias with a LLM agreement/disagreement analysis in Appendix.Sec. C to gain further insights. In particular, the disagreement analysis helps identify instances where it is easy or difficult for judges to reach a consensus, reflecting the complexity of the judgment task. Based on our findings regarding the answer quality gap and LLM agreement/disagreement, this study offers practical implications for designing evaluator benchmarks that control the difficulty of judging tasks. Our findings show that the hardest-to-judge instances are those where (1) LLMs frequently disagree with each other, (2) comparable win rates and minimal quality gaps exist between the two candidate models, and (3) significant position bias is exhibited by the majority of judges.
One qualitative analysis example of a difficult-to-judge instance comes from MTBench question 120, which involved a math problem where both candidate models, Vicuna-13b-v1.3 and WizardLM-30b, provided incorrect solutions. The problem was: (1) Given $f(x)=4x^{3}-9x-14$, find the value of $f(2)$, and (2) Find $x$ such that $f(x)=0$. Vicuna-13b-v1.3 incorrectly applied the quadratic formula to solve the cubic equation, while WizardLM-30b used an enumeration method but failed to identify the correct answer. This resulted in a 4{A}-3{B}-5{C} choice pattern among the twelve judges, illustrating a typical case of significant disagreement.
We observed that some judges exhibited hallucinations, making false positive judgments or ignoring the reference answer. Others preferred either correctness (choosing {C} as a tie) or methodological soundness (choosing {B}, favoring WizardLM-30b). In the swapped order setting, 9 out of 12 judges displayed position bias. This highlights the challenge of judging solutions with minimal quality gaps, leading to greater position bias and subjectivity.
The key takeaway is that reducing hard-to-evaluate, highly subjective instances can improve benchmark design. Likewise, minimizing trivial cases where judges easily agree would enhance the benchmark’s robustness. Clearer and more specific evaluation criteria could also help resolve ambiguity in difficult instances.
@@ 5 Conclusion
In conclusion, this paper proposes a systematic, scalable framework for evaluating the position bias of LLM judges in pairwise comparative scenarios in terms of repetition stability, position consistency, and preference fairness. Through comprehensive evaluations of 12 judges on two benchmarks across 22 tasks with over 100,000 evaluation instances, we observe significant variations of position bias across judges and tasks. We also discover that position bias is weakly length-dependent but significantly impacted by the answer quality gap. Our findings enhance the understanding of position bias, paving the way for more effective mitigation strategies and more robust and reliable evaluation systems.
@@
Appendix
Appendix A Contribution & Limitation
Contributions
Despite the limited number of models, tasks, and judging types, our proposed systematic framework is agnostic to these criteria and can be used with any LLM judge. Our work is particularly valuable for validating judges in terms of repetition stability and then evaluating them through position consistency and preference fairness. We also explore the factors influencing position bias and their quantitative impacts, such as the answer quality gap and lengths of prompt components. Particularly, our study is the first to establish a direct connection between quantified answer quality disparities and position bias, highlighting the urgent need for a comprehensive understanding of how position bias affects LLM judges. Additionally, the mutual agreement and disagreement analysis of LLM judges are insightful for identifying the difficulty of evaluation instances and beneficial for benchmark assessment and designs. All findings from our study enhance the understanding of the position bias of LLM-as-a-Judge, paving the way for more effective mitigation strategies and more robust and reliable evaluation systems.
Social Benefits Our comprehensive understanding of position bias in LLM-as-a-Judge provides multiple benefits to the community. First, it enhances the reliability and trustworthiness of LLM judges in terms of repetition stability validation. Our focus and investigation on preference fairness inspire future works to improve the fairness of LLM evaluators simultaneously with consistency and reliability. Moreover, our emphasized trade-offs between position consistency, fairness, and practical applicability when selecting task-specific LLM judges provide crucial judge-selection guidance for future work. Additionally, the baseline comparison analysis across judges and tasks offers reasonable judge model recommendations for more reliable, consistent, and fair evaluation results. Besides, the LLM agreement/disagreement analysis in Appendix.Sec.C 5 provides an applicable tool to identify the difficulty of judging certain instances. Together with findings from the significant impact of the answer quality gap on position bias, our study benefits future benchmark designs and evaluations and lays the groundwork for potential mitigation strategies of position bias. Overall, our study contributes to a more reliable, unbiased evaluation system that benefits the application of LLM evaluators across diverse fields, such as healthcare (Chen et al., 2024c), instruction following and prompting (Zeng et al., 2023; Chua et al., 2024), multimodal assessment (Chen et al., 2024a), and recommender systems (Hou et al., 2024).
Strength
Our framework of systematically understanding position bias has the following strengths: (1) scalability and ease of use - our proposed framework is scalable to various judges, models, tasks, and judging types. It is also straightforward and convenient to implement, facilitating a more comprehensive evaluation of LLM judges for future works. (2) comprehensive experiments - we investigate the position bias of 12 LLM judges on 2 benchmarks across 22 tasks and around 40 answer-generating models in terms of repetition stability, position consistency, and preference fairness, resulting in more than 100,000 evaluation instances; the exploration of the factors impacting position bias from Judge-level, Model-level, and Task-level and the findings from LLM agreement analysis provide additional insights to enhance the understanding of position bias. (3) accurate measurement - we formally define crucial metrics ($RS$ and $PF$) for evaluating LLM judges’ position bias that are underexplored by prior works and offer detailed explanations for their rationality and rigorousness.
Limitations
Due to computational limits, our study can be potentially extended to provide more comprehensive insights. (1) We only study the pairwise comparative assessment of LLM-as-a-Judge for its best consistency, ease of use, and scalability, but more types of judging can be explored, such as pairwise scoring and listwise ranking. (2) We only study 12 up-to-date commercial LLM judges for systematic investigations. Llama 7b-13b models that we choose as representatives of open-source models of similar sizes, face extreme difficulty in judging the instances explored in this study. Due to limited resources and considering practicality, we did not experiment with open-source models with larger sizes. However, further exploration may include analysis of these state-of-the-art open-source and fine-tuned judge models using our proposed framework. (3) We only study position bias on two benchmarks, MTBench and DevBench, which is still limited despite 22 tasks and 40 answer-generating models. Also, the number of questions for each task is limited. Further studies can be extended to more benchmarks to have a broader understanding of position bias across benchmarks, tasks, and questions. (4) Due to data accessibility, we are unable to investigate Judge-level factors in terms of model parameter sizes and architectures, which we expect will have a significant impact on position bias. Alternatively, we consider the family property to account for Judge-level factors by grouping models with similar properties (e.g., approximate parameter size) together. We also investigate variances of position bias across individual judges. However, future exploration on these model properties, especially those available on open-source models, will provide more insights.
Future Work
Besides addressing the limitations, the following future work, built on findings from our study, could provide more insights: (1)prompt setting - in this study, we apply the exact default prompt settings of MTBench and DevBench. However, the positional order and prompt style of not just the model-generated responses but also system prompt components (e.g., agent role assignment, mode of judging, direct mention of not making biases) may also have an impact on the extent to which LLM judges make position bias. (2) open-source insights: Further investigations of open-source and fine-tuned models on easier-to-evaluate benchmark may provide more insights into the training stages or model structures that are affecting or causing the position bias. Then our framework offers better control on the other impactful factors such as the answer quality gap to make these further investigations more accurate.
@@ B.1 LLM-as-a-Judge
In recent years, Large Language Models (LLMs) have emerged as a transformative technology, garnering global attention and stimulating substantial research into their applications. For evaluative tasks, particularly subjective ones, human assessment is considered the gold standard due to its comprehensive and open-ended nature (Zeng et al., 2023). However, it lacks scalability and reproducibility (Karpinska et al., 2021).
As a result, LLMs have increasingly been used as substitutes for human evaluators across various Natural Language Generation (NLG) domains and tasks (Zheng et al., 2024b; Li et al., 2024; Chiang & Lee, 2023b; Liusie et al., 2024; Chen et al., 2023; Fu et al., 2023; Wang et al., 2023a; Kocmi & Federmann, 2023b; Ji et al., 2023; Liu et al., 2024; Zeng et al., 2023; Liu et al., 2023; Shen et al., 2023; Kocmi & Federmann, 2023a; Chiang & Lee, 2023a; Karpinska et al., 2021; Dubois et al., 2024; Kim et al., 2023; Liang et al., 2023; Chang et al., 2023; Zhong et al., 2023; Chia et al., 2023; Deb & Das, 2023; Ye et al., 2024; Min et al., 2023; Li et al., 2023d; c; b; Lu et al., 2024), including open-ended story generation (Chiang & Lee, 2023a), adversarial attacks (Chiang & Lee, 2023b), summarization (Karpinska et al., 2021), machine translation (Kocmi & Federmann, 2023a), and instruction following (Zeng et al., 2023).
These LLM evaluators, known as LLM-as-a-Judge, have attracted significant interest within both academic and industrial circles (Zheng et al., 2024b; a; Pezeshkpour & Hruschka, 2023; Chiang & Lee, 2023b; Shen et al., 2023; Liusie et al., 2024; Qin et al., 2024; Chen et al., 2023; Huang et al., 2024; Zhu et al., 2023b; Raina et al., 2024; Chen et al., 2024b; Wang et al., 2023b; Liu et al., 2024; Zeng et al., 2023; Khan et al., 2024; Kim et al., 2023; Li et al., 2023b; d; Shi et al., 2024; Li et al., 2023c; Chen et al., 2024c; Chua et al., 2024; Chen et al., 2024a). As LLMs have made content generation significantly easier, the volume of generated responses has increased, making it impractical to rely solely on human evaluation. Therefore, cost-effective LLM judges are needed to assess these responses efficiently. LLM-as-a-Judge is typically employed in Q&A evaluation tasks, where the LLM judge is prompted to evaluate the quality of responses, usually generated by models, answering the questions.
In many of these tasks, LLM judges have shown a high level of agreement with human evaluators (Zheng et al., 2024b; Li et al., 2024; Zhu et al., 2023b), yet in some tasks, they are less effective, largely due to inherent biases (Shen et al., 2023; Chiang & Lee, 2023b; Zheng et al., 2024a). Even in cases where agreement is high, judgments may still suffer from biases.
When employing LLM-as-a-Judge, various types of judging are available, which can be categorized either by the scale or comparative method. From the scale perspective, LLM-as-a-Judge can involve pointwise, listwise, or pairwise assessment (Qin et al., 2024). By comparative method, it can be either score-based or relation-based (Li et al., 2023d).
For example, pointwise scoring (Kim et al., 2023) lets the LLM judge score the response of one model to a question at a time based on some evaluating metrics. Pairwise/Listwise Scoring (Karpinska et al., 2021; Zhu et al., 2023b; Chen et al., 2024a; Zheng et al., 2024b) prompts the LLM judges to score a pair/list of model-generated answers. Listwise Ranking is another relation-based assessment that, instead of giving a score, requires the LLM judge to rank a list of responses following some specified order (e.g., from best to worst). Pairwise comparative assessment (Karpinska et al., 2021; Chen et al., 2024a; Zheng et al., 2024b), on the other hand, asks the LLM judge to select the superior response between a given pair, usually conducted in a double-blind manner: the generating model remains unknown to the judge, and the judging model remains anonymous to the answer generator.
Except for pointwise evaluation, all forms of LLM-as-a-Judge suffer—or are susceptible to suffering—from position bias due to the intrinsic nature of position and comparison within the prompt structure.
@@ B.2 Position Bias
Previous studies have discovered and investigated multiple types of biases, such as position bias (Zheng et al., 2024b; Li et al., 2024; Chen et al., 2024a; 2023; Chia et al., 2023; Chiang & Lee, 2023b; Khan et al., 2024; Zhu et al., 2023b; Zhong et al., 2023; Zheng et al., 2024a; Deb & Das, 2023; Zeng et al., 2023; Wang et al., 2023c; Qin et al., 2024; Wang et al., 2023b; Shi et al., 2024; Liusie et al., 2024; Liu et al., 2024; Kim et al., 2023; Karpinska et al., 2021; Hou et al., 2024; Kocmi & Federmann, 2023b; Pezeshkpour & Hruschka, 2023; Shen et al., 2023; Chiang & Lee, 2023a; Kocmi & Federmann, 2023a; Chen et al., 2024b; Li et al., 2023c; d; Raina et al., 2024), verbosity/length bias (Zheng et al., 2024b; Liu et al., 2024; Kim et al., 2023), self-enhancement bias (Li et al., 2023c; Zheng et al., 2024b), selection bias (Zheng et al., 2024a; Pezeshkpour & Hruschka, 2023), and contextual bias (Liu et al., 2024; Zhou et al., 2024).
Among these, position bias stands out as particularly significant, permeating a wide array of tasks and affecting judge models, including open-source (Li et al., 2023b; Chen et al., 2024c), proprietary commercial ones (OpenAI, 2023; Anthropic, 2024; Gemini Team, 2024), and fine-tuned models (Huang et al., 2024; Wang et al., 2023c; Zhu et al., 2023b; Li et al., 2023b; Kim et al., 2023).
To clarify, the aforementioned and following position bias refers to the concept within the context of LLM-as-a-Judge, meaning that LLM judge tends to favor responses based on their position in the prompt rather than their content. For example, in a pairwise comparative assessment scenario, if the LLM judge consistently selects the first response as superior even after switching the order of the two responses (same position, but different content), then a position bias occurs. Some other studies also use the term position bias (Ko et al., 2020; Aslanyan & Porwal, 2019; Blunch, 1984; Raghubir & Valenzuela, 2006), but in this research, our interest lies solely in the position bias specific to LLM-as-a-Judge.
Position bias is arguably the most prevalent and impactful bias among all. Chua et al. (2024) notes that their Bias-Augmented Consistency Training (BCT), an unsupervised fine-tuning scheme designed to promote consistent reasoning across prompts with and without biasing features, improves Chain-of-Thought (Wei et al., 2022) performance over self-training controls for all biases except position bias. Furthermore, Khan et al. Khan et al. (2024) point out that LLM judges are less confident when exhibiting position bias, and addressing this bias is highly complex due to varying confidence levels across judges and tasks.
Moreover, there is ongoing debate over whether selection bias originates from position bias. Pezeshkpour & Hruschka (2023) argue that LLMs are sensitive to the ordering of options in Multiple Choice Questions (MCQs), confirming that position bias contributes to this sensitivity. Raina et al. (2024) contest this view, asserting that selection bias stems less from position bias and more from token bias, which represents an inherent challenge for LLMs and contributes to poor robustness.
All of these findings underscore the critical importance of addressing the issue of position bias when employing LLM-as-a-Judge.
B.3 Deal with Position Bias
Intuitively, position bias emerges because LLMs are sensitive to changes, especially positional changes, in prompts (Zheng et al., 2024b; Raina et al., 2024; Pezeshkpour & Hruschka, 2023; Zhao et al., 2021; Zhu et al., 2023a). Also, LLM judges are vulnerable to attacks (Raina et al., 2024; Chen et al., 2024b; Shi et al., 2024).
Parse data with position bias There are many ways to deal with the position bias. The naive way is to exclude the inconsistent judgments(i.e. if the LLM judge gives a positionally biased judgment on a pair or list of model responses) (Zheng et al., 2024b; Chen et al., 2024b; Wang et al., 2023b; Li et al., 2023b). While this ensures consistent and reliable remaining judgments, it does not resolve the fundamental problem. Moreover, if the LLM judge is highly biased, this method discards valuable evaluation instances and information, making it an ineffective and somewhat desperate measure.
To take the positionally inconsistent evaluations into account, one may either take an average for scoring-based judging (Zheng et al., 2024b; Raina et al., 2024; Li et al., 2023b; Wang et al., 2023b) or regard inconsistency as a half-win or tie for relation-based judging (Zheng et al., 2024b; Li et al., 2023c) after swapping the order of model responses in the prompt. For instance, in a pairwise scoring scenario, if model A receives a score of 8 when put in the first position and 4 when put in the second position, the overall score for it compared to model B would be 6; in the pairwise comparative case, if model A wins when put in the first position but then lose when put in the second position, it counts as a tie or half-win for both model A and model B. The latter way of swapping + tie is proposed because intuitively the position bias is more likely to occur when the model responses get evaluated share a similar quality in terms of the evaluating metric. Our quantitative study also verifies this intuition, evidenced by the fact the LLM judge’s positional consistency (the percentage of positionally consistent pairs of judgment) is positively proportional to the answer quality gap between the model responses.
To save the expense of running the experiments more than once using swapped order, in practice, many of these studies also support a random-shuffle option in their code settings such that the baseline model for comparison does not remain in a fixed position.
Solution attempts Due to the significance of position bias, more sophisticated and advanced approaches have emerged to solve position bias as well, including bootstrapping (Hou et al., 2024), split-and-merge (Li et al., 2023d), and multi-agent discussion (Li et al., 2023c; Khan et al., 2024). However, these methods are either costly and time-consuming (e.g., multi-agent discussion and review to reach agreement) or ineffective.
Furthermore,Liu et al. (2024) suggests that existing calibration techniques designed to reduce bias, including context decomposition (Li et al., 2023d), order permutation (Wang et al., 2023b; Zheng et al., 2024a), ensembling (Li et al., 2023a), and batch calibration (Zhou et al., 2024), are insufficient to align LLM evaluators, even with supervised data.
Thus, position bias is a pervasive, significantly impactful, and challenging problem to solve.
Understand before solving We observe that the existing methods maybe ineffective or not enough satisfactory because there is a lack of understanding of the position bias. Although a variety of studies have researched this type of bias in LLM-as-a-Judge, a comprehensive understanding of what factors affect the position bias remains underexplored. In other words, without clarity on the key factors and their quantitative impact on position bias, the efficacy of current and future methods remains uncertain.
For instance, Li et al. (2023d) proposed a PORTIA approach that addresses the position bias to a large extent, receiving a 80.99% Fixed Coverage (the percentage of positionally inconsistent original assessments that are later corrected by PORTIA) for GPT-4 in a relation-based evaluation on 8 MTBench answer pairs, improving consistency from 93.44 % to 97.03%. However, the extraordinarily high consistency of the original evaluation may imply that the choice of answer pairs may be biased in terms of answer quality gap, a factor that our study proves to be significantly impactful. In other words, if the quality between the two model responses differs considerably, the LLM judges will make little position bias during judgment and hence is easy to calibrate. This example illustrates how lacking a comprehensive understanding of the factors affecting position bias can lead to overestimating or misjudging the effectiveness of methods proposed to resolve it and improve the performance of LLM-as-a-Judge.
Missing analysis for position bias Besides, the positional preference direction aspect of position bias remains underexplored. Positional preference refers to the specific positions an LLM judge favors when position bias is evident. For instance, if an LLM judge consistently favors the responses that appear first in the prompt, it exhibits a preference for the first position. Our study denotes primacy-preferred and recency-preferred to describe biases toward the first and second positions, respectively. We integrate the measurement by formally defining positional fairness score to quantify the extent of a judge model’s preference towards certain positions.
While prior research includes some positional preference results (Zheng et al., 2024b; Wang et al., 2023b; Zhu et al., 2023b; Li et al., 2023c; Liusie et al., 2024), none have conducted a detailed analysis. They focus primarily on positional consistency in relation to position bias, while we regard positional preference as an equally crucial component that needs to be examined, understood, and enhanced in future work.
Therefore, we propose that mitigating the position bias of LLM-as-a-Judge requires a simultaneous improvement in both positional consistency and positional fairness. A positionally consistent (high positional consistency) and positionally fair (nearly equal preference on primacy and recency when position bias occur) LLM judge is undoubtedly preferable. However, a trade-off between consistency and fairness/preference often occurs in practice. Previous studies that have provided results on positional preference have all shown that even state-of-the-art (SOTA) LLM judges struggle with both positional inconsistency and clear positional preferences.
Apart from that, previous studies have yet to examine how stable LLM judges are across repetitions. MLLM-as-a-Judge (Chen et al., 2024a) conducts repeated experiments but then takes an average/mode of the judgment only to make the evaluations more robust and reliable, without a focus on the repetitively stable judgments. In comparison, we investigate the repetition stability to determine whether position bias is solely influenced by the prompt’s positional information and the judge’s intrinsic properties or can be partly attributed to judgment randomness. If the judgments exhibit randomness over repetitions, then the position bias overlaps with repetition bias, complicating the issue and potentially invalidating previous findings.
B.4 Summary of Prior Work
To summarize, LLM-as-a-Judge has a large potential in alternating human judges on a wide range of tasks, in particular subjective evaluation, owing to its cost-effectiveness, high agreement with human judgments, reproducibility, and scalability. However, it suffers from various biases, most notably position bias, which is prevalent across different evaluation tasks, models, and judgment types. Not only is it a type of bias that significantly hinders the improvement and promotion of applying LLM-as-a-Judge, but it is also difficult to solve due to its complexity and lack of understanding in the community.
Existing solutions proposed to address this issue may be ineffective or uncertain in efficacy due to the unclear impact and nature of influencing factors. Additionally, whether the position bias is an essential bias or mere random variation has not been validated from prior works. The positional preference side, reflecting the positional fairness of the LLM judge, also requires comprehensive analysis. In a nutshell, a comprehensive understanding of the position bias of LLM-as-a-Judge is crucial to validate existing and future approaches to address this important problem.

### v7

@@ 4 Main Findings
Position Bias is weakly Length-dependent We investigate the impact of three different lengths on the position bias of LLM judges: the length of the question (task input length), the solution length of candidate models (task output length), and the length of the entire prompt (prompt length). By stepwise regression, we discovered that average task output length is only significant in predicting $PF$, adding a minimal change in AIC as shown in Appendix Table. 4. In other words, there is a very weak relationship between the lengths of prompt components and position bias.
LLM Agreement Analysis: We complement our investigation of position bias with an agreement and disagreement analysis among LLM judges. Rather than focusing exclusively on overall consistency or fairness, we examine how judges converge and diverge in their assessments of individual instances. Agreement analysis quantifies the percentage of instances where two judges mutually agree on the outcome. Disagreement analysis counts the number of choices deviating from the mode for each instance among all judges. This further complies a “distribution of disagreement” across the dataset.
Our findings reveal that, despite exhibiting similar overall $PC$ and $PF$ scores, judges vary significantly in their judgments on individual instance. Disagreement analysis, in particular, highlights instances where consensus is either easily or difficultly achieved, reflecting the inherent complexity of the judgment task. For example, as shown in Fig. 3, more than half of the dataset can be considered relatively easy to judge, as over 75% of all 12 judges agree with each other on these instances. Conversely, fewer than 4% of instances represent the likely hard-to-judge cases where a majority of judges fail to reach consensus.
Figure 3: Distribution of disagreement on MTBench. The y-axis indicates the proportion of the dataset where the level of disagreement among judges does not exceed a specific threshold.
Based on our observations of answer quality gaps and LLM agreement/disagreement patterns, this study offers practical insights for designing evaluator benchmarks that account for the varying difficulty levels of judgment tasks. Specifically, the most challenging instances to evaluate are characterized by: (1) frequent disagreements among LLM judges, (2) closely matched win rates and minimal quality gaps among candidate models, and (3) significant position bias exhibited by the majority of judges. Further discussions and analyses can be found in Appendix Sec. C.
@@ 5 Conclusion
In conclusion, this paper provides an in-depth analysis of position bias in LLM judges, a critical challenge in automated evaluation. Using metrics such as repetition stability, position consistency, and preference fairness, we identify significant variations in position bias across judges and tasks, consistent across pairwise and list-wise comparison settings. Our findings show that position bias is weakly influenced by prompt length but strongly impacted by the quality gap between solutions. Furthermore, agreement and disagreement analysis highlights variability in judgment reliability, providing valuable insights for refining benchmarks. This study enhances understanding of position bias and contributes to the development of fairer and more reliable LLM evaluation frameworks.
6 Limitations
Despite proposing scalable metrics and investigating key factors influencing position bias, our study has several limitations. First, due to computational constraints, we evaluated only twelve commercial LLM judges for pairwise settings and four for list-wise settings, limited to 3-candidate lists, using two benchmarks. Open-source models, such as Llama 7B-13B, struggled to provide reliable judgments, restricting their inclusion. Additionally, while we tested original benchmark prompt templates, exploring variations in prompting techniques could provide further insights. Future work could expand this study to include more models, tasks, prompting strategies, and larger list-wise candidate pools to validate and generalize our findings.
Second, data accessibility limitations prevented a direct analysis of Judge-level factors like architecture and parameter size. Instead, we approximated these factors by grouping models by family properties. Open-source models with accessible architectural details could allow deeper analysis of these influences. Additionally, our analyses rely on completed judgments, such as measuring the answer quality gap and task difficulty after evaluations. Future work could explore methods to estimate or control these factors pre-judgment, reducing computational costs and enabling proactive mitigation strategies.
Lastly, our focus was on understanding position bias rather than mitigating it. While our findings provide a foundation for effective mitigation, further research is needed to address issues like maintaining consistency and fairness when answer quality gaps are minimal, where position bias is most pronounced. Multivariate analyses exploring interactions between factors like prompt length, task complexity, and answer quality gaps could also yield deeper insights and enhance mitigation approaches.
@@ 6 Limitations
Appendix A Reproducibility
Our experiments were conducted primarily using API access, with a total cost of approximately 3,000 USD.
The code repository for reproducing our results is available at: https://anonymous.4open.science/r/Position-Bias-Analyzer-Demo-F7E3
@@ B.1 LLM-as-a-Judge
Large Language Models (LLMs) have become a transformative tool in automating evaluative tasks, offering scalability and reproducibility advantages over human assessments. The methodology of using LLMs as evaluators ("LLM-as-a-Judge") has been widely used for tasks such as open-ended story generation (Chiang and Lee, 2023a), adversarial attacks (Chiang and Lee, 2023b), summarization (Karpinska et al., 2021), machine translation (Kocmi and Federmann, 2023), and instruction following (Zeng et al., 2023), where models are tasked with scoring or ranking outputs. Despite their potential, inherent biases—particularly position bias—pose significant challenges to their reliability and fairness, even in the most effective pairwise comparative settings.
@@ B.2 Position Bias
However, position bias is a complex problem that is challenging to deal with or solve. The naive way is to exclude the position inconsistent judgments, which does not solve the fundamental issue and would likely result in data sparsity when position bias is frequently exhibited. Therefore, researchers proposed ‘inconsistency-as-a-tie’ for both models in pairwise comparative settings to consider all judgments for further analysis (Zheng et al., 2024b; Li et al., 2023c). This approach, while practically useful for evaluations, does not mitigate position bias.
Given the significance of position bias, researchers have developed more sophisticated approaches to mitigate it, including bootstrapping (Hou et al., 2024), split-and-merge techniques (Li et al., 2023d), and multi-agent discussions (Li et al., 2023c; Khan et al., 2024). While these methods demonstrate potential, they are often costly, time-consuming, or insufficient. For example, multi-agent discussions and iterative reviews require extensive computational resources and human-like deliberation, while simpler techniques, such as order permutation (Wang et al., 2023; Zheng et al., 2024a) and batch calibration (Zhou et al., 2024), have limited effectiveness in addressing bias holistically. Furthermore, calibration methods like ensembling (Li et al., 2023a) or context decomposition (Li et al., 2023d) are often constrained by the inadequacy of supervised data to align LLM evaluators effectively (Liu et al., 2024). This highlights that addressing this crucial bias is complex and challenging.
The limitations of existing approaches underscore a critical gap in our understanding of position bias itself. A lack of clarity on the factors influencing this bias and their quantitative impacts hampers the development of effective solutions. For instance, while the PORTIA approach (Li et al., 2023d) demonstrates significant improvements in consistency for GPT-4 (from 93.44% to 97.03%), this success may be overestimated due to the selection of tasks with large answer quality gaps. As our study reveals, such gaps inherently reduce the likelihood of position bias, making these cases easier to calibrate. This example underscores the necessity of first comprehensively understanding the underlying factors contributing to position bias before proposing or refining mitigation strategies. Without this foundational understanding, the efficacy of current and future methods remains uncertain and potentially misjudged.
B.3 Research Gaps
To enhance the understanding of position bias, our study builds on and extends prior work through the following dimensions.
Repetition Stability: Previous studies on LLM-as-a-Judge often implicitly assume that single-instance judgments are sufficient to reflect the evaluating capabilities of judge models, thereby attributing observed biases entirely to systematic tendencies rather than random variations. These studies typically proceed directly to analyses on position bias or others without such verification. However, if judgments vary by chance or lack consistency across repetitions, the reliability and accuracy of downstream analyses are compromised. To avoid this issue, Chen et al. (2024) repeated evaluations multiple times and adopted the mode judgment as the representative, an approach that, while effective, is computationally expensive. Addressing this gap, our work introduces a method or metric for systematically evaluating the stability of LLM judgments across repetitions. This ensures that observed biases are systematic rather than artifacts of variability. By performing this validation on a sample dataset for cost-efficiency, researchers can establish a robust foundation for subsequent evaluations and analyses, ensuring the reliability of their conclusions.
Preference Fairness: Prior studies on position bias often focus on the proportion of position consistent judgments, referred to as “position consistency”. Only a few have explored the preference directions of LLM evaluators in addition to consistency. However, traditional measurements either overly penalize highly consistent judges (Li et al., 2023c; Liusie et al., 2024) or rely on metrics whose sensitivity is highly dependent on dataset size (Zheng et al., 2024b; Zhu et al., 2023). To address these limitations, we proposed a refined metric, preference fairness ($PF$), which integrates the strengths of both approaches. This single, comprehensive metric ensures sensitivity across datasets, remains independent of the number of questions and tasks, and represents a significant improvement over previous metrics.
Answer Quality Gap: Intuitively, candidate solutions with significant quality disparities are easier to judge, likely resulting in reduced position bias. This suggests that the quality gap among answers is a potentially significant factor influencing position bias. However, this factor has been largely overlooked in prior works, both in evaluations and in the development of mitigation strategies, underscoring our limited understanding of position bias. While Wang et al. (2023) explored the answer quality gap in score-based LLM-as-a-Judge evaluations using judgment scores, quantifying this variable in comparative settings (pairwise or list-wise) remains an open research problem. To address this gap, we propose the use of win rates (Zheng et al., 2024b) of candidate solutions against an expected baseline to measure the answer quality gap. This approach offers a practical quantitative metric for subsequent analysis and deepens our understanding of the factors driving position bias.
List-wise Evaluations: Although LLM-as-a-Judge has been categorized into pointwise, pairwise, and list-wise paradigms (Qin et al., 2024), most research has focused on pairwise comparative assessments due to their demonstrated effectiveness (Zheng et al., 2024b; Liusie et al., 2024). However, this leaves a significant gap in measuring and understanding the position bias in list-wise settings, where LLMs are prompted to evaluate three or more candidate solutions simultaneously. With advancements in model capabilities, list-wise evaluations present a potentially more cost-effective alternative to pairwise approaches, requiring fewer comparisons while maintaining consistency and fairness. Our work bridges this gap by extending position bias assessments from pairwise to list-wise settings, enabling a more comprehensive evaluation of LLM-as-a-Judge across varied configurations.
These identified gaps - spanning from the validation of judgments to the evaluation of list-wise settings - highlight the need for a more comprehensive framework to evaluate and understand position bias. By addressing these dimensions collectively, our work provides a foundation for a deeper understanding of the consistency, fairness, and reliability of LLM evaluators.